The classified balance sheet for a company reported current assets of $1,624,125, total liabilities of $810,540, common stock of $1,110,000, and retained earnings of $141,260. The current ratio was 2.5. Which of the following statements is not correct?

1. Total Assets are $2,061,800.

2. Total Stockholders’ equity is $1,251,260.

3. Noncurrent liabilities are $141,260.

4The amount of current assets is 2.5 times the amount of current liabilities.
Answer #1

Total stockholders equity=Common stock+Retained earnings

=(1,110,000+141260)=$1,251,260

Total assets=Total stockholders equity+Total liabilities

=$1,251,260+$810,540=$2,061,800

Current ratio=Current assets/Current liabilities

Hence Current assets=2.5*Current liabilities.

Current liabilities=(1,624,125/2.5)=$649,650

Non current liabilities=Total liabilities-Current liabilities

=810,540-649,650=$160,890.

Hence statement 3 is not correct.